A delivery truck carrying 75 canvas of rice is traveling at 10.0 m/s. The mass of the truck and one ca van of rice are 3000 kg and 50 kg, respectively. The driver of the truck has a mass of 85 kg.

(a)   what is the magnitude of the momentum of the truck with the driver and the canvas of rice?

(b)  At a certain grocery, the driver unloaded eight canvas. What is the new momentum-of the truck with the driver and the canvas of rice if the driver maintains the same speed?


1
Expert's answer
2021-03-23T08:14:01-0400

a)


"p=mv\\p=(3000+75\cdot50)(10)=67500\frac{kgm}{s}"

b)


"p'=(3000+(75-8)\cdot50)(10)=63500\frac{kgm}{s}"